# A Kinetic Model of Solar Wind Acceleration Driven by Ambipolar Electric Potential and Velocity-Space Diffusion

### v1

@@ Abstract
Results. The effect of velocity-space diffusion is found to reduce the temperature anisotropy and impact the solar wind acceleration while maintaining the electric potential unchanged. The approach described in this work allows for adjusting the diffusion to effectively reduce or increase the solar wind acceleration. This potentially provides a way to account for the unexplained acceleration within the kinetic exospheric framework. It is found that, even without diffusion, the model is able to reproduce the anticorrelation between the electric potential and the solar wind terminal velocity found by PSP. This suggests that the electric potential might still be of major importance in explaining the solar wind acceleration.
@@ 1 Introduction
The electrons and protons that are blown away from the Sun through interplanetary space are the main constituents of a continuous outflow of particles referred to as the solar wind. The exact mechanism that allows the solar wind to reach observed velocities averaging between 300 and 800 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ remains one of the most significant open questions in space physics (e.g. Rouillard et al., 2021). Since the first in situ solar wind observations in 1959, neither magnetohydrodynamic (MHD) nor kinetic models have been able to self-consistently explain the solar wind acceleration. In particular, MHD models are not able to capture the origin of the fast solar wind with realistic coronal temperatures and often require additional heating mechanisms (Parker 1965; Leer et al. 1982; Hansteen & Velli 2012). Although some kinetic approaches can provide an explanation for the presence of peculiar observed velocity distribution functions (VDFs) of solar wind particles and propose potential mechanisms of solar wind acceleration (e.g. Hollweg & Isenberg 2002; Vocks et al. 2005; Seough et al. 2015), they generally lack any global electric field-driven component of the acceleration. In this paper, we investigate this acceleration by improving a specific class of kinetic models known as exospheric models, which are fundamentally based on considering the electric field-driven component as the primary acceleration mechanism. A comprehensive review of these models has been provided by Fahr & Shizgal (1983).
@@
The first exospheric models were developed for planetary exospheres and were later adapted for the solar wind (Chamberlain 1960). These models describe the radial evolution of the particle VDFs by neglecting collisions between particles. The first model of the solar wind that included a consistent “ambipolar” electric field was developed by Lemaire & Scherer (1971) in a collisionless regime considered above the solar surface, at a given altitude called the “exobase”, which was chosen at 6 solar radii ($R_{s}$). This electric field is called ambipolar in the sense that it satisfies the quasi-neutrality and zero-current conditions between protons and electrons, observed to be preserved in the solar wind (e.g. Feldman et al., 1975). The challenging aspect of these models was to determine the radial evolution of this ambipolar electric field self-consistently with the exospheric approach, especially in the case of non-monotonic proton potential (typically found when assuming an exobase below 6 $R_{s}$, which better describes the solar wind case), as noted by Jockers (1970). Inspired by the latter, a generalized method to approximate the solution for this self-consistent electric potential was later developed by Lamy et al. (2003).
@@
The use of Kappa (also known as Lorentzian) VDFs for electrons, in contrast to Maxwellian VDFs for protons, constitutes one of the major breakthroughs that was first introduced in an exospheric model by Pierrard & Lemaire (1996) and applied to the solar wind by Maksimovic et al. (1997a). This approach was motivated by the need to account for the so-called suprathermal electrons observed extensively in the solar wind (e.g. Lin 1980; Pilipp et al. 1987; Pierrard et al. 2016; Abraham et al. 2022; Salem et al. 2023). Such suprathermal electron populations are better described by a Kappa distribution function rather than a single Maxwellian distribution, as shown, for instance, by Maksimovic et al. (1997b) and more recently by Zheng et al. (2024). These electrons are high-energy particles that do not follow the classical single Maxwellian distribution fitted to the lower-energy part (called the core) of the observed distribution. They form the halo, which extends in all directions, and the electron strahl, a more concentrated component observed in the anti-sunward direction (e.g.  Feldman et al. 1975; Pilipp et al. 1987; Pierrard et al. 2001b). Evidence has been found that indicates the presence of the strahl even within the solar corona, as observed with Parker Solar Probe (PSP) for both slow and fast winds (Halekas et al. 2022). The mathematical description of Kappa distributions required for exospheric models was first introduced by Pierrard & Lemaire (1996) and generalized for non-monotonic potentials by Lamy et al. (2003). Within the scope of this model, using a Kappa distribution for electrons increases the number of electrons with high speeds in the anti-sunward direction, which effectively increases the electric potential difference between any two radial distances (see Pierrard & Lazar 2010 for a review). Therefore, this allows more protons to reach the maximum of the non-monotonic proton potential (usually within 1 to 5 $R_{s}$), which in turn accelerates the wind to realistically high bulk speeds without requiring unrealistically high temperatures. Following the same approach, Zouganelis et al. (2004) showed that similar results could also be obtained with a sum of two Maxwellians or a sum of Maxwellian and Kappa distributions (for core and strahl), which would not limit the significance of the strahl to the power law factor $\kappa$.
@@ 2 Review of the Solar Wind Kinetic Exospheric Model
$\frac{\mathcal{E}_{p}}{k_{b}T_{0p}}\equiv\frac{m_{p}v^{2}}{2k_{b}T_{0p}}+\frac{m_{p}\Delta\phi+Z_{p}e\Delta V}{k_{b}T_{0p}}=\frac{v_{\parallel}^{2}+v_{\perp}^{2}}{w_{p}^{2}}+\psi_{p}(r),$ (7)
@@
Here, $\kappa$ is the parameter associated with a power law decreasing Kappa VDF, instead of an exponentially decreasing Maxwellian VDF (i.e. for $\kappa\rightarrow\infty$ the Maxwellian distribution is retrieved); $w_{e}=\sqrt{(2\kappa-3)k_{b}T_{0e}/(\kappa m_{e})}$ is the electron thermal velocity, $T_{0e}$ is the electron temperature at $r_{0}$, $m_{e}$ is the electron mass, $\Gamma$ is the Gamma function, $N_{e}$ is a factor of normalization (defined in Section 3) that allows for retrieving the specified electron density at $r_{0}$ by integrating the electron VDF in velocity space, and $\mathcal{E}_{e}/[(\kappa-3/2)k_{b}T_{0e}]$ is the normalized electron energy so that
$\begin{split}\frac{\mathcal{E}_{e}}{(\kappa-3/2)k_{b}T_{0e}}&\equiv\frac{m_{e}v^{2}}{2(\kappa-3/2)k_{b}T_{0e}}+\frac{m_{e}\Delta\phi+Z_{e}e\Delta V}{(\kappa-3/2)k_{b}T_{0e}}\\ &\equiv\frac{(v_{\parallel}^{2}+v_{\perp}^{2})}{\kappa w_{e}^{2}}+\frac{\psi_{e}(r)}{\kappa},\end{split}$ (10)
@@
Table 1: Input parameters of the kinetic exospheric model used for the VDFs shown in Fig. 1
@@
These dashed lines delimit the regions of velocity space labeled with “A”, “B”, and “C” in Fig. 1. The A region corresponds to the particles that do not have sufficient energy to overcome the potential maximum; these are called “ballistic” particles. The B region corresponds to particles that have enough energy to overcome this potential maximum and are therefore called “escaping” particles. The C region corresponds to trapped particles that lack sufficient energy to overcome the potential maximum and cannot reach $r_{0}$. Instead, they are perpetually reflected because of magnetic moment conservation. In principle, such particles cannot originate directly from $r_{0}$. However, as generally assumed (e.g. Maksimovic et al. 1997a; Lamy et al. 2003; Zouganelis et al. 2004), our model still accounts for them because this region is expected to eventually become saturated with particles: any particle entering this region of velocity space (whether through rare collisions or interactions with plasma waves) will remain there indefinitely unless scattered out. Note that in our example there is no region for trapped protons (Fig. 1d), but this is not always the case. In fact, with the right parameters, trapped particles could be obtained close to $r_{0}$ (see, e.g., Lamy et al. 2003).
@@
The solutions for $\psi_{p}(r_{m})$, $\psi_{e}(r_{\infty})$, and $r_{m}$ are determined through an iterative method, employing these equations and the requirement that $\psi_{p}(r)$ must be unique and continuous (Jockers 1970, Lamy et al. 2003). This enables the calculation of the potentials $\psi_{p}(r)$ and $\psi_{e}(r)$, which are subsequently applied to compute macroscopic quantities using moments equations, provided by Lamy et al. (2003) and Pierrard & Lemaire (1996), involving the integrals of the VDFs in different regions of velocity space (i.e. A, B, or C) for protons and electrons, respectively.
@@ 3 Exospheric Approach on Diffusion
The pressures are those corresponding to the D region of the electron VDF in Fig. 2, where $N_{e,\textbf{D}}$ is the normalization factor of the electron VDF so that the density of this region is always equal to $n_{e,\textbf{D}}(r)$ as described by Equation (31). This is inferred by noticing that
@@
With this relatively simple approach, the influence of an arbitrary diffusion on the radial evolution of the density, bulk velocity, and temperature can be calculated with our kinetic exospheric model.
@@ 4 Diffusive Kinetic Exospheric Model: Results
The radial evolution of the electron densities is shown in Fig. 4 for the parameters of Table 1 and the same diffusion as previously considered ($\zeta=0.1$). As mentioned in the previous section, the total density is unaffected by the diffusion. It can be seen that the amount of diffused electrons experiences a rapid growth with radial distances, exceeding the number of escaping electrons (i.e. within the B region) just above the exobase (at 1.2 $R_{s}$). The diffused particles even exceed the ballistic ones at around 32 $R_{s}$ (not shown in this figure). This occurs because the trapped particles (that is, those within the C region) rapidly dominate the total electron density. Since the number of diffused particles is determined by the total density, the diffused electrons subsequently outweigh the ballistic ones at large distances.
@@ 5 Exospheric Perspective on Solar Wind Acceleration
For the kinetic exospheric model, the radial evolution of the self-consistent electric potential is shown in Fig. 5 for four different sets of parameters given by Table 2. The SSW1 solid black line is a slow solar wind solution found using the same test parameters as in Table 1, which was introduced to illustrate the effect of diffusion. Note that the electric potential is independent of the diffusion, since it is found by using the quasi-neutrality with an unchanged total density. It is commonly understood that exospheric models predict a faster solar wind accompanied by a larger electric potential at any radial distance, to account for the enhanced acceleration of the solar wind. This contradiction with the observed anticorrelation mentioned above holds true only if we consider that the acceleration is entirely determined by the $\kappa$ parameter. As illustrated in Fig. 5, this is evident from the comparison between SSW1 and FSW1, where reducing $\kappa$ from 6 to 3 results in a corresponding increase in the terminal wind velocity from 320 to 640 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$. However, when considering the same change in $\kappa$ for SSW2 and FSW2, a carefully selected combination of exobase levels and initial temperatures (see Table 2) produces a crossing point between the two electric potentials (at around 13 $R_{s}$) while still maintaining the increase in terminal wind velocity from 280 (for SSW2) to 520 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ (for FSW2). In this way, the anticorrelation between the electric potential and the solar wind terminal velocity is retrieved above 13 $R_{s}$ with the kinetic exospheric model. This indicates that the kinetic exospheric model may be better suited than previously thought for explaining the fast solar wind through the electric field, especially if the crossing point between the electric potentials of fast and slow winds occurs below 13 $R_{s}$. Although PSP observations do not appear to indicate the presence of such an intersection, or at least not above 13 $R_{s}$, the consistency of our modeled electric potentials compared with observations suggests that further investigation of the most recent PSP data is needed to completely rule out this possibility.
@@
It is important to note that the implementation of the diffusion mechanism does not inherently require that particles in region D do not contribute to the flux. With this consideration, the contribution of diffused particles to the flux could potentially exceed its depletion within the strahl caused by the diffusion itself. In that case, a greater diffusion would imply a greater acceleration of the solar wind. Therefore, diffusion may serve as a basis for an acceleration mechanism that could be incorporated into the kinetic exospheric model using the approach outlined in this paper. On top of a possible flux contribution of the halo, this diffusion framework could also be generalized to include a proper sunward deficit (not just an empty region) as well as an amplification of the strahl population, thus again introducing an additional acceleration within the kinetic exospheric model.
The observed suprathermal electrons, included in our existing kinetic exospheric model, constitute an important property of the strahl as they may have significant implications on the solar wind acceleration. According to Maksimovic et al. (2021), spectral line analyzes seem to indicate that suprathermal populations are present in the corona. However, in situ data from PSP presented by Halekas et al. (2022) suggest that no significant strahl suprathermal electrons are observed near the Sun. This observation implies that Kappa electron VDFs may not be the most likely explanation for the fast solar wind. However, our results, consistent with the findings of Zouganelis et al. (2004), indicate that the key factor influencing the solar wind’s terminal velocity is not solely the suprathermal component, but rather the relative abundance of particles within the strahl compared to the deficit in the sunward direction.

### v2

@@ Abstract
Results. The effect of velocity-space diffusion is found to reduce the temperature anisotropy and impact the solar wind acceleration while maintaining the electric potential unchanged. The approach described in this work allows for adjusting the diffusion to effectively reduce or increase the solar wind acceleration. It is found that, even without diffusion, the model is able to reproduce the anticorrelation between the electric potential and the solar wind terminal velocity found by PSP. This suggests that the electric potential might still be of major importance in explaining the solar wind acceleration.
@@ 1 Introduction
The electrons and protons that are blown away from the Sun through interplanetary space are the main constituents of a continuous outflow of particles referred to as the solar wind. The exact mechanism that allows the solar wind to reach observed velocities averaging between 300 and 800 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ remains one of the most significant open questions in space physics (e.g., Rouillard et al. 2021). Since the first in situ solar wind observations in 1959, neither magnetohydrodynamic (MHD) nor kinetic models have been able to self-consistently explain the solar wind acceleration. In particular, MHD models are not able to capture the origin of the fast solar wind with realistic coronal temperatures and often require additional heating mechanisms (Parker 1965; Leer et al. 1982; Hansteen & Velli 2012). Although some kinetic approaches can provide an explanation for the presence of peculiar observed velocity distribution functions (VDFs) of solar wind particles and propose potential mechanisms of solar wind acceleration (e.g., Hollweg & Isenberg 2002; Vocks et al. 2005; Seough et al. 2015), they generally lack any global electric field-driven component of the acceleration. In this paper, we investigate this acceleration by improving a specific class of kinetic models known as exospheric models, which are fundamentally based on considering the electric field-driven component as the primary acceleration mechanism. A comprehensive review of these models has been provided by Fahr & Shizgal (1983).
@@
The first exospheric models were developed for planetary exospheres and were later adapted for the solar wind (Chamberlain 1960). These models describe the radial evolution of the particle VDFs by neglecting collisions between particles. The first model of the solar wind that included a consistent “ambipolar” electric field was developed by Lemaire & Scherer (1971) in a collisionless regime considered above the solar surface, at a given altitude called the “exobase”, which was chosen at 6 solar radii ($R_{s}$). This electric field is called ambipolar in the sense that it satisfies the quasi-neutrality and zero-current conditions between protons and electrons, observed to be preserved in the solar wind (e.g., Feldman et al. 1975). The challenging aspect of these models was to determine the radial evolution of this ambipolar electric field self-consistently with the exospheric approach, especially in the case of non-monotonic proton potential (typically found when assuming an exobase below 6 $R_{s}$, which better describes the solar wind case), as noted by Jockers (1970). Inspired by the latter, a generalized method to approximate the solution for this self-consistent electric potential was later developed by Lamy et al. (2003).
@@
The use of Kappa (also known as Lorentzian) VDFs for electrons, in contrast to Maxwellian VDFs for protons, constitutes one of the major breakthroughs that was first introduced in an exospheric model by Pierrard & Lemaire (1996) and applied to the solar wind by Maksimovic et al. (1997a). This approach was motivated by the need to account for the so-called suprathermal electrons observed extensively in the solar wind (e.g., Lin 1980; Pilipp et al. 1987; Pierrard et al. 2016; Abraham et al. 2022; Salem et al. 2023). Such suprathermal electron populations are better described by a Kappa distribution function rather than a single Maxwellian distribution, as shown, for instance, by Maksimovic et al. (1997b) and more recently by Zheng et al. (2024). These electrons are high-energy particles that do not follow the classical single Maxwellian distribution fitted to the lower-energy part (called the core) of the observed distribution. These suprathermal distributions are typical for the halo, which extends in all directions, and the electron strahl, a more concentrated component observed in the anti-sunward direction (e.g., Feldman et al. 1975; Pilipp et al. 1987; Pierrard et al. 2001b). The mathematical description of Kappa distributions required for exospheric models was first introduced by Pierrard & Lemaire (1996) and generalized for non-monotonic potentials by Lamy et al. (2003). Within the scope of this model, using a Kappa distribution for electrons increases the number of electrons with high speeds in the anti-sunward direction, which effectively increases the electric potential difference between any two radial distances (see Pierrard & Lazar 2010 for a review). Therefore, this allows more protons to reach the maximum of the non-monotonic proton potential (usually within 1 to 5 $R_{s}$), which in turn accelerates the wind to realistically high bulk speeds without requiring unrealistically high temperatures. Following the same approach, Zouganelis et al. (2004) showed that similar results could also be obtained with a sum of two Maxwellians or a sum of Maxwellian and Kappa distributions (for core and strahl), which would not limit the significance of the strahl to the power law factor $\kappa$.
@@ 2 Review of the solar wind kinetic exospheric model
$\frac{\mathcal{E}_{p}}{k_{b}T_{0p}}\equiv\frac{m_{p}v^{2}}{2k_{b}T_{0p}}+\frac{m_{p}\Delta\phi+Z_{p}e\Delta V}{k_{b}T_{0p}}=\frac{v_{\parallel p}^{2}+v_{\perp p}^{2}}{w_{p}^{2}}+\psi_{p}(r),$ (7)
@@
Here, $\kappa$ is the parameter associated with a power law decreasing Kappa VDF, instead of an exponentially decreasing Maxwellian VDF (i.e., for $\kappa\rightarrow\infty$ the Maxwellian distribution is retrieved); $w_{e}=\sqrt{(2\kappa-3)k_{b}T_{0e}/(\kappa m_{e})}$ is the electron thermal velocity, $T_{0e}$ is the electron temperature at $r_{0}$, $m_{e}$ is the electron mass, $\Gamma$ is the Gamma function, $N_{e}$ is a factor of normalization (defined in Section 3) that allows for retrieving the specified electron density at $r_{0}$ by integrating the electron VDF in velocity space, and $\mathcal{E}_{e}/[(\kappa-3/2)k_{b}T_{0e}]$ is the normalized electron energy so that
$\begin{split}\frac{\mathcal{E}_{e}}{(\kappa-3/2)k_{b}T_{0e}}&\equiv\frac{m_{e}v^{2}}{2(\kappa-3/2)k_{b}T_{0e}}+\frac{m_{e}\Delta\phi+Z_{e}e\Delta V}{(\kappa-3/2)k_{b}T_{0e}}\\ &\equiv\frac{(v_{\parallel e}^{2}+v_{\perp e}^{2})}{\kappa w_{e}^{2}}+\frac{\psi_{e}(r)}{\kappa},\end{split}$ (10)
@@
Table 1: Input parameters of the kinetic exospheric model used for the VDFs shown in Fig. 1.
@@
These dashed lines delimit the regions of velocity space labeled with “A”, “B”, and “C” in Fig. 1. The A region corresponds to the particles that do not have sufficient energy to overcome the potential maximum; these are called “ballistic” particles. The B region corresponds to particles that have enough energy to overcome this potential maximum and are therefore called “escaping” particles. The C region corresponds to trapped particles that lack sufficient energy to overcome the potential maximum and cannot reach $r_{0}$. Instead, they are perpetually reflected because of magnetic moment conservation. In principle, such particles cannot originate directly from $r_{0}$. However, as generally assumed (e.g., Maksimovic et al. 1997a; Lamy et al. 2003; Zouganelis et al. 2004), our model still accounts for them because this region is expected to eventually become saturated with particles: any particle entering this region of velocity space (whether through rare collisions or interactions with plasma waves) will remain there indefinitely unless scattered out. Therefore, we consider that the VDF of region C is also described by Equations (6) and (9) for protons and electrons respectively, ensuring VDF continuity between ballistic and trapped particles. Note that in our example there is no region for trapped protons (Fig. 1d), but this is not always the case. In fact, with the right parameters, trapped particles could be obtained close to $r_{0}$ (see, e.g., Lamy et al. 2003).
@@
The solutions for $\psi_{p}(r_{m})$, $\psi_{e}(r_{\infty})$, and $r_{m}$ are determined through an iterative method, employing these equations and the requirement that $\psi_{p}(r)$ must be unique and continuous (Jockers 1970, Lamy et al. 2003). This enables the calculation of the potentials $\psi_{p}(r)$ and $\psi_{e}(r)$, which are subsequently applied to compute macroscopic quantities using moments equations, provided by Lamy et al. (2003) and Pierrard & Lemaire (1996), involving the integrals of the VDFs in different regions of velocity space (i.e., A, B, or C) for protons and electrons, respectively.
@@ 3 Exospheric approach on diffusion
The pressures are those corresponding to the D region of the electron VDF in Fig. 2, where $N_{e,\textbf{D}}$ is the normalization factor of the electron VDF (replacing $N_{e}$ in Equation (9)) so that the density of this region is always equal to $n_{e,\textbf{D}}(r)$ as described by Equation (31). This is inferred by noticing that
@@
With this relatively simple approach, the influence of diffusion on the radial evolution of the density, bulk velocity, and temperature can be calculated with our kinetic exospheric model.
@@ 4 Diffusive kinetic exospheric model: results
The radial evolution of the electron densities is shown in Fig. 4 for the parameters of Table 1 and the same diffusion as previously considered ($\zeta=0.1$). As mentioned in the previous section, the total density is unaffected by the diffusion. It can be seen that the amount of diffused electrons experiences a rapid growth with radial distances, exceeding the number of escaping electrons (i.e., within the B region) just above the exobase (at 1.2 $R_{s}$). The diffused particles even exceed the ballistic ones at around 32 $R_{s}$ (not shown in this figure). This occurs because the trapped particles (that is, those within the C region) rapidly dominate the total electron density. Since the number of diffused particles is determined by the total density, the diffused electrons subsequently outweigh the ballistic ones at large distances.
@@ 5 Exospheric perspective on solar wind acceleration
For the kinetic exospheric model, the radial evolution of the self-consistent electric potential is shown in Fig. 5 for four different sets of parameters given by Table 2. The SSW1 solid black line is a slow solar wind solution found using the same test parameters as in Table 1, which was introduced to illustrate the effect of diffusion. Note that the electric potential is independent of the diffusion, since it is found by using the quasi-neutrality with an unchanged total density. It is commonly understood that exospheric models predict a faster solar wind accompanied by a larger electric potential at any radial distance, to account for the enhanced acceleration of the solar wind. This contradiction with the observed anticorrelation mentioned above holds true only if we consider that the acceleration is entirely determined by the $\kappa$ parameter. As illustrated in Fig. 5, this is evident from the comparison between SSW1 and FSW1, where reducing $\kappa$ from 6 to 3 results in a corresponding increase in the terminal wind velocity from 320 to 640 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$. However, when considering the same change in $\kappa$ for SSW2 and FSW2, a carefully selected combination of exobase levels and initial temperatures (see Table 2) produces a crossing point between the two electric potentials (at around 13 $R_{s}$) while still maintaining the increase in terminal wind velocity from 280 (for SSW2) to 520 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ (for FSW2). These parameters were chosen because they approximate the observed behavior to some extent, producing a crossing point at low altitude while maintaining realistic values (as described in the next paragraph). However, a wide range of parameter sets can reproduce a crossing point, provided they meet two key conditions: (1) a sufficiently large difference between proton and electron temperatures (compared to the difference in kappa values) in fast wind relative to slow wind, and (2) a significantly higher exobase level for slow wind. In this way, the anticorrelation between the electric potential and the solar wind terminal velocity is retrieved above 13 $R_{s}$ with the kinetic exospheric model. This indicates that the kinetic exospheric model may be better suited than previously thought for explaining the fast solar wind through the electric field, especially if the crossing point between the electric potentials of fast and slow winds occurs below 13 $R_{s}$. Although PSP observations do not appear to indicate the presence of such an intersection, or at least not above 13 $R_{s}$, the consistency of our modeled electric potentials compared with observations suggests that further investigation of the most recent PSP data is needed to completely rule out this possibility.
@@
It is important to note that the implementation of the diffusion mechanism does not inherently require that particles in region D do not contribute to the flux. For that reason, an additional acceleration mechanism could be taken into account within the current diffusion framework by considering that the transfer of core particles (combined ballistic and trapped particles of region A and C) to higher energies contributes to the flux. In practice, this transfer would either increase the number of escaping particles, enhance the anti-sunward portion of the halo (compared to the sunward portion), or both. Another refinement for better alignment with observations would be to account for diffusion creating a proper sunward deficit, rather than just an empty region, along with its associated negative flux contribution. The key idea is that the sum of all newly introduced flux contributions from diffusion (of the core) could ultimately result in a net increase in total flux, thereby providing an additional acceleration mechanism within the kinetic exospheric model.
One limitation to consider is that the escaping particles (region C) defined within our current kinetic exospheric approach cannot be the sole source of halo electrons. A quantitative analysis of our exospheric model results (neglecting diffusion) indicates that the relative density of escaping electrons, compared to the total density, is approximately 3% at 65 $R_{s}$, increasing to about 4% at 215 $R_{s}$, with minimal variations across the parameter sets used in this study (SSW1, FSW1, SSW2, and FSW2). However, observations by Štverák et al. (2009) suggest that the combined halo and strahl population should constitute roughly 7.5% of the total electron density over the same radial distances. This discrepancy implies that escaping particles alone cannot account for the observed halo electron population within our current exospheric model framework.
Taking this into account, we introduced a diffusion that transfers particles from the strahl to the halo, which improves the representation of the combined halo and strahl populations, yielding a relative density of approximately 8% from 65 to 215 $R_{s}$, in better agreement with observations. Importantly, the primary objective of this study is to propose a modification to the exospheric framework that incorporates diffusion in the simplest possible manner, as this represents the first exospheric model of its kind. However, existing diffusion models (e.g., Verscharen et al. 2019; Jeong et al. 2020, 2022), supported by observations of the strahl and halo populations (e.g., Štverák et al. 2008, 2009), indicate that diffusion primarily occurs from the strahl to the halo. This suggests that further refinement of our kinetic exospheric model is necessary to ensure consistency with established models and observations.
The observed suprathermal electrons, included in our existing kinetic exospheric model, constitute an important property of the strahl as they may have significant implications on the solar wind acceleration. According to Effenberger & Jeffrey (2021), spectral line analyzes seem to indicate that suprathermal populations are present in the corona. However, in situ data from PSP presented by Halekas et al. (2022) suggest that no significant strahl suprathermal electrons are observed near the Sun. This observation implies that Kappa electron VDFs may not be the most likely explanation for the fast solar wind. However, our results, consistent with the findings of Zouganelis et al. (2004), indicate that the key factor influencing the solar wind’s terminal velocity is not solely the suprathermal component, but rather the relative abundance of particles within the strahl compared to the deficit in the sunward direction.